# Chirped DFT-s-OFDM: A new single-carrier waveform with enhanced LMMSE noise suppression

### v1

@@ III-B Output SNR Analysis Using LMMSE Equalizer
By letting $\mathbf{H}_{\rm{eff}}=\mathbf{F}_{N}\mathbf{C}^{H}\textbf{H}_{\mathrm{t}}\mathbf{C}\mathbf{F}_{N}^{H}\mathbf{P}\mathbf{F}_{M}$, (5) is rewritten as
@@
$\hat{\mathbf{x}}_{\rm{LMMSE}}$ $=\mathbf{H}_{\rm{eff}}^{H}\left(\mathbf{H}_{\rm{eff}}\mathbf{H}_{\rm{eff}}^{H}+\mathbf{I}_{N}\sigma^{2}\right)^{-1}\mathbf{y}$
$=\underbrace{\mathbf{H}_{\rm{eff}}^{H}\left(\mathbf{H}_{\rm{eff}}\mathbf{H}_{\rm{eff}}^{H}+\mathbf{I}_{N}\sigma^{2}\right)^{-1}\mathbf{H}_{\rm{eff}}\mathbf{x}}_{\textrm{Desired signal}}$
$+\underbrace{\mathbf{H}_{\rm{eff}}^{H}\left(\mathbf{H}_{\rm{eff}}\mathbf{H}_{\rm{eff}}^{H}+\mathbf{I}_{N}\sigma^{2}\right)^{-1}\mathbf{z}}_{\rm{Noise}}.$ (13)
@@
Output SNR $=\frac{\mathbb{E}\lVert\mathbf{x}^{H}\mathbf{H}_{\rm{eff}}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{H}_{\rm{eff}}\mathbf{x}\rVert}{\mathbb{E}\lVert\mathbf{z}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{z}\rVert}$
$=\frac{\mathbb{E}\{\rm{Tr}\{\mathbf{x}\mathbf{x}^{H}\mathbf{H}_{\rm{eff}}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{H}_{\rm{eff}}\}\}}{\mathbb{E}\{\rm{Tr}\{\mathbf{z}\mathbf{z}^{H}\mathbf{G}^{H}\mathbf{G}\}\}}$
$=\frac{\mathbb{E}\{\rm{Tr}\{\mathbf{H}_{\rm{eff}}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{H}_{\rm{eff}}\}\}}{\mathbb{E}\{\rm{Tr}\{\sigma^{2}\mathbf{G}^{H}\mathbf{G}\}\}},$ (14)
where $\mathbf{G}=\mathbf{H}_{\rm{eff}}^{H}\left(\mathbf{H}_{\rm{eff}}\mathbf{H}_{\rm{eff}}^{H}+\mathbf{I}_{N}\sigma^{2}\right)^{-1}$, $\mathbb{E}(\cdot)$ and $\rm{Tr}(\cdot)$ indicate the expectation and trace operation.
Figs. 4a, 4b, and 4c depict the output signal power, output noise power, and output SNR using LMMSE equalizer of proposed chirped DFT-s-OFDM, in comparison to existing waveforms. According to Fig. 4a, the proposed waveform has slightly higher output signal power than DFT-s-OFDM, AFDM, and OTFS, while OFDM has deep fades with signal power reduced significantly. In Fig. 4b, the proposed chirped DFT-s-OFDM has the lowest output noise power. DFT-s-OFDM, AFDM, and OTFS have similar noise power while OFDM has two spikes with very high noise power. In Fig. 4c, the proposed chirped DFT-s-OFDM has highest output SNR and the output SNR remains nearly identical at each symbol. The output SNRs of DFT-s-OFDM, AFDM, and OTFS are close and roughly half of that of chirped DFT-s-OFDM. Though OFDM has good output SNR at some symbols, its deep fades greatly degrade the overall performance. In conclusion, the proposed chirped DFT-s-OFDM can enhance noise suppression of LMMSE equalization, and thus, would contribute to enhanced communication performance.
@@ IV Simulation Results
The communication performances of proposed chirped DFT-s-OFDM are studied and compared with the existing waveforms, e.g., DFT-s-OFDM, AFDM, OTFS, and OFDM. Unless otherwise stated, the values of simulation parameters are set in the following. The IFFT size of chirped DFT-s-OFDM, DFT-s-OFDM, AFDM, and OFDM is $N=256$. The DFT size of chirped DFT-s-OFDM and DFT-s-OFDM is $M=64$, leading to the DFT spreading factor of $\rm{SF}=4$. The chirp rate is $c=1/N$. The carrier frequency and subcarrier spacing are $f_{\rm{c}}=4\,\rm{GHz}$ and $\Delta f=15\,\rm{KHz}$. The $3$-path equal-gain channel is assumed. The velocity is $v=500\,\rm{km/h}$, corresponding to Doppler frequency of $f_{\rm{max}}=2\,\rm{KHz}$. For OTFS, the numbers of delay grids and Doppler grids are $M_{\rm{OTFS}}=16$ and $N_{\rm{OTFS}}=16$. Quadrature phase shift keying (QPSK) modulation and LMMSE equalizer are adopted.
Figure 5: PAPRs of proposed chirped DFT-s-OFDM and other waveforms.
@@
Fig. 5 shows the complementary cumulative distribution function (CCDF) of PAPRs of proposed chirped DFT-s-OFDM and existing waveforms. The proposed waveform by multiplying chirp signal with DFT-s-OFDM preserves the low PAPR of DFT-s-OFDM. Their PAPR is much lower than that of multi-carrier waveforms, e.g., AFDM, OTFS, and OFDM.
@@
Figure 6: a) BER and b) output SNR of proposed chirped DFT-s-OFDM, in comparison with other waveforms.
@@
Fig. 6 presents the BER and output SNR of proposed chirped DFT-s-OFDM, compared to DFT-s-OFDM, AFDM, OTFS, and OFDM. Thanks to enhanced noise suppression discussed in Subsections II-B and III-B, the proposed chirped DFT-s-OFDM outperforms the existing waveforms in terms of BER and output SNR, with around $4\,\rm{dB}$ $\rm{SNR}$ gain over AFDM, OTFS, and DFT-s-OFDM. Though OFDM has output SNR that is as good as that of OTFS and AFDM, it has deep fades as shown in Fig. 4. Meanwhile, it cannot exploit frequency diversity, and thus, its BER is worst in Fig. 6a).
@@
Fig. 7a) studies the impact of DFT spreading factor on BER of DFT-s-OFDM and proposed chirped DFT-s-OFDM. The BER of DFT-s-OFDM would not obviously enhance with DFT spreading factor. In contrast, for the proposed chirped DFT-s-OFDM, there is an obvious BER improvement especially when DFT spreading factor increases from $2$ to $4$. This is because the increase of DFT spreading factor tends to enhance noise suppression as can be seen in Fig. 7b). As discussed in Subsection II-B, the enhanced noise suppression results from full band transmission and symbols retransmission enabled by chirping and DFT precoding, respectively.
@@
Figure 7: Impact of DFT spreading factor on a) BER and b) output noise power of DFT-s-OFDM and proposed chirped DFT-s-OFDM.
@@ V Conclusion
In this correspondence, a new single-carrier waveform called chirped DFT-s-OFDM has been proposed for 6G communications. The new waveform by chirping DFT-s-OFDM in the time domain maintains the low PAPR of DFT-s-OFDM. Thanks to the full band transmission and symbols retransmission enabled by chirping and DFT precoding, the proposed waveform enables enhanced noise suppression of LMMSE equalization. It presents higher output SNR and lower BER than the state-of-the-art waveforms, including DFT-s-OFDM, AFDM, OTFS, and OFDM. In addition, the PEP analysis and simulation results confirm that the proposed waveform can achieve full frequency diversity.

### v2

@@ III-B Output SNR Analysis Using LMMSE Equalizer
This subsection aims to analyze output SNR of proposed chirped DFT-s-OFDM using LMMSE equalizer. By letting $\mathbf{H}_{\rm{eff}}=\mathbf{F}_{N}\mathbf{C}^{H}\textbf{H}_{\mathrm{t}}\mathbf{C}\mathbf{F}_{N}^{H}\mathbf{P}\mathbf{F}_{M}$, (4) is rewritten as
@@
$\hat{\mathbf{x}}_{\rm{LMMSE}}=\mathbf{H}_{\rm{eff}}^{H}\left(\mathbf{H}_{\rm{eff}}\mathbf{H}_{\rm{eff}}^{H}+\mathbf{I}_{N}\sigma^{2}\right)^{-1}\mathbf{y}=\mathbf{G}\mathbf{H}_{\rm{eff}}\mathbf{x}+\mathbf{G}\mathbf{z},$ (13)
with $\mathbf{G}=\mathbf{H}_{\rm{eff}}^{H}\left(\mathbf{H}_{\rm{eff}}\mathbf{H}_{\rm{eff}}^{H}+\mathbf{I}_{N}\sigma^{2}\right)^{-1}$. Note that intersymbol interference (ISI) has been mitigated to some extent through LMMSE equalizer. According to (13), without considering the impact of noise, the signal estimate is expressed as $\mathbf{G}\mathbf{H}_{\rm{eff}}\mathbf{x}$. The impact of ISI can be viewed from the non-diagonal elements of $\mathbf{G}\mathbf{H}_{\rm{eff}}$, which are small and negligible.
@@
Output $\textrm{Output\,SNR}=\frac{\mathbb{E}\{\mathbf{x}^{H}\mathbf{H}_{\rm{eff}}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{H}_{\rm{eff}}\mathbf{x}\}}{\mathbb{E}\{\mathbf{z}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{z}\}},$ (14)
where $\mathbb{E}(\cdot)$ indicates the expectation operation. Due to the properties of trace operation, the following can be obtained: $\mathbf{x}^{H}\mathbf{H}_{\rm{eff}}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{H}_{\rm{eff}}\mathbf{x}=\rm{Tr}\{\mathbf{x}^{H}\mathbf{H}_{\rm{eff}}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{H}_{\rm{eff}}\mathbf{x}\}=\rm{Tr}\{\mathbf{x}\mathbf{x}^{H}\mathbf{H}_{\rm{eff}}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{H}_{\rm{eff}}\}$ and $\mathbf{z}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{z}=\rm{Tr}\{\mathbf{z}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{z}\}=\rm{Tr}\{\mathbf{z}\mathbf{z}^{H}\mathbf{G}^{H}\mathbf{G}\}$. (14) can be rewritten as
Output $\textrm{Output\,SNR}=\frac{\mathbb{E}\{\rm{Tr}\{\mathbf{x}\mathbf{x}^{H}\mathbf{H}_{\rm{eff}}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{H}_{\rm{eff}}\}\}}{\mathbb{E}\{\rm{Tr}\{\mathbf{z}\mathbf{z}^{H}\mathbf{G}^{H}\mathbf{G}\}\}}$
$=\frac{\rm{Tr}\{\mathbb{E}\{\mathbf{x}\mathbf{x}^{H}\}\mathbb{E}\{\mathbf{H}_{\rm{eff}}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{H}_{\rm{eff}}\}\}}{\rm{Tr}\{\mathbb{E}\{\mathbf{z}\mathbf{z}^{H}\}\mathbb{E}\{\mathbf{G}^{H}\mathbf{G}\}\}}=\frac{\mathbb{E}\{\rm{Tr}\{\mathbf{H}_{\rm{eff}}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{H}_{\rm{eff}}\}\}}{\mathbb{E}\{\rm{Tr}\{\sigma^{2}\mathbf{G}\mathbf{G}^{H}\}\}},$ (15)
with $\mathbb{E}\{\mathbf{x}\mathbf{x}^{H}\}=\mathbf{I}_{M}$, $\mathbb{E}\{\mathbf{z}\mathbf{z}^{H}\}=\sigma^{2}\mathbf{I}_{N}$, and $\rm{Tr}\{\mathbf{G}^{H}\mathbf{G}\}=\rm{Tr}\{\mathbf{G}\mathbf{G}^{H}\}$. As a result, the output signal power and output noise power are the sum of diagonal elements of $\mathbf{H}_{\rm{eff}}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{H}_{\rm{eff}}$ and $\sigma^{2}\mathbf{G}\mathbf{G}^{H}$, respectively. The $m$-th diagonal element of $\mathbf{H}_{\rm{eff}}^{H}\mathbf{G}^{H}\mathbf{G}\mathbf{H}_{\rm{eff}}$ and $\sigma^{2}\mathbf{G}\mathbf{G}^{H}$ correspond to the output signal power and output noise power at $m$-th symbol, and they are shown in Figs. 3a and 3b, respectively. The ratio of Fig. 3a and Fig. 3b yields the output SNR with respect to symbol index $m$ and is shown in Fig. 3c. Following the similar derivations, the output signal power, output noise power, and output SNR are plotted for other waveforms in Fig. 3.
First, according to Fig. 3, unlike (chirped) DFT-s-OFDM, AFDM, and OTFS, the output signal power, output noise power, and output SNR of OFDM differ for some symbols. This is because OFDM symbols are modulated in the frequency domain and different symbol would experience different fading due to multipath channel. For other waveforms, each symbol is transmitted across nearly the entire frequency band, enhancing their resilience to frequency-selective fading. Second, the proposed chirped DFT-s-OFDM produces lower output noise power and higher output SNR than DFT-s-OFDM. This is due to the inclusion of chirping, facilitating full-band transmission. For example, in Fig. 1b, with $\mathrm{SF}=4$, $2$ unknown data symbols can be estimated from the received data symbols at $8$ subcarriers for chirped DFT-s-OFDM rather than $2$ subcarriers for DFT-s-OFDM. There are more observations than unknowns and the redundant information from more equations can mitigate the impact of noise. Third, the chirped DFT-s-OFDM outperforms AFDM and OTFS in terms of output noise power and output SNR. This is because the introduction of DFT precoding in chirped DFT-s-OFDM spreads data symbols in the time domain. For example, in Fig. 1b, the data symbols are transmitted two times or four times over time, and the retransmission could reduce the impact of noise.
@@ IV Simulation Results
The performance of chirped DFT-s-OFDM is evaluated and compared with the existing waveforms in single-user uplink MAC, including DFT-s-OFDM, AFDM, OTFS, and OFDM. To ensure a fair comparison, different waveforms are compared with identical bit to noise ratio Eb/N0. The energy per (active) subcarrier for (chirped) DFT-s-OFDM remains the same to that for OFDM, AFDM, and OTFS. Unless otherwise stated, the values of simulation parameters are set as follows. The IFFT size of (chirped) DFT-s-OFDM, AFDM, and OFDM is $N=256$. The DFT size of (chirped) DFT-s-OFDM is $M=64$, leading to the DFT spreading of $\rm{SF}=4$. The chirp rate is $c_{\rm{r}}=1/N$. The carrier frequency and subcarrier spacing are $f_{\rm{c}}=4\,\rm{GHz}$ and $\Delta f=15\,\rm{KHz}$. The $3$-path equal-gain channel ($L=3$) is assumed. The velocity is $v=500\,\rm{km/h}$, corresponding to Doppler frequency of $f_{\rm{max}}=2\,\rm{KHz}$. For each Monte Carlo simulation, the values of Doppler shifts are randomly generated from $-f_{\rm{max}}$ to $f_{\rm{max}}$. For OTFS, the numbers of delay grids and Doppler grids are $M_{\rm{OTFS}}=16$ and $N_{\rm{OTFS}}=16$. Quadrature phase shift keying (QPSK) modulation and LMMSE equalizer are adopted.
@@
Fig. 4a shows the complementary cumulative distribution function (CCDF) of PAPRs of proposed chirped DFT-s-OFDM and existing waveforms. The proposed waveform by multiplying chirp signal with DFT-s-OFDM preserves the lower PAPR of DFT-s-OFDM over AFDM/OFDM and OTFS. Note that (chirped) DFT-s-OFDM in Fig. 4a has a PAPR with $0\,\rm{dB}$. This is because DFT-s-OFDM symbol after interleaved subcarrier mapping is the original un-modulated symbols with reduced amplitude and its PAPR remains identical to that of un-modulated symbols [2, 3, 4]. When phase shift keying (PSK) modulation is employed, a PAPR of $0\,\rm{dB}$ can thus be achieved.
@@
Figure 4: a) PAPR, b), BER , and c) output SNR of proposed chirped DFT-s-OFDM, in comparison with existing waveforms.
@@
Figs. 4b and 4c present the BER and output SNR of proposed chirped DFT-s-OFDM, compared to DFT-s-OFDM, AFDM, OTFS, and OFDM. Thanks to enhanced noise suppression discussed in Subsections II-B and III-B, the proposed chirped DFT-s-OFDM outperforms the existing waveforms in terms of BER and output SNR, with around $4\,\rm{dB}$ $\rm{SNR}$ gain over AFDM, OTFS, and DFT-s-OFDM. Though OFDM has output SNR that is as good as that of OTFS and AFDM, it has deep fades as shown in Fig. 3. Meanwhile, it cannot exploit frequency diversity, and thus, its BER is worst in Fig. 4b).
@@
Fig. 5a) studies the impact of DFT spreading factor $\mathrm{SF}$ on BER of (chirped) DFT-s-OFDM. The BER of DFT-s-OFDM would not obviously enhance with $\mathrm{SF}$. For the proposed chirped DFT-s-OFDM, there is an obvious BER improvement especially when $\mathrm{SF}$ increases from $2$ to $4$. This is because the increase of $\mathrm{SF}$ tends to enhance noise suppression as can be seen in Fig. 5b). As discussed in Subsection II-B, the enhanced noise suppression results from full band transmission and symbols retransmission enabled by chirping and DFT precoding, respectively. After $\mathrm{SF}=4$, the performance enhancement of chirped DFT-s-OFDM is not obvious. This is because the increase of $\mathrm{SF}$ would reduce signal amplitude by $1/\sqrt{\mathrm{SF}}$ [2, 3, 4]. The amplitude reduction is significant as $0.35$ and $0.25$ for $\mathrm{SF}=8$ and $\mathrm{SF}=16$, respectively. This amplitude reduction would result in performance degradation, cancelling out the additional benefits gained from additional repeated transmission. Thus, when $\mathrm{SF}$ further exceeds, there is no substantial BER improvement.
@@
Figure 5: Impact of DFT spreading factor $\mathrm{SF}$ on a) BER and b) output noise power of (chirped) DFT-s-OFDM.
Note that, in the previous results, (chirped) DFT-s-OFDM with $\mathrm{SF}>1$ exhibits a different spectral efficiency compared to AFDM and OTFS using all subcarriers. Fig. 6a shows the BERs of different modulation waveforms under the same bandwidth efficiency with $\mathrm{SF}=4$. AFDM and OTFS use a similar data mapping scheme as (chirped) DFT-s-OFDM, with one-fourth of the interleaved subcarriers (for AFDM) or delay grids (for OTFS) allocated to a specific user for data transmission. Note that the unused subcarriers or delay grids can be assigned to other users for data transmission. Compared to Fig. 4b, the BERs of AFDM and OTFS in Fig. 6a are lower and are nearly identical to that of chirped DFT-s-OFDM. This implies that the superior BER of chirped DFT-s-OFDM over AFDM and OTFS in Figs. 3 and 4 mainly results from DFT precoding. Though AFDM, OTFS, and chirped DFT-s-OFDM exhibit similar BER in Fig. 6a, the first two produce higher PAPR than the third. In practical implementation, when the input signal is too high, the power amplifier’s output would reach its maximum limit and is unable to increase further, resulting in clipping [12, 13]. When the clipping ratio is set to $1$ [13], the BERs of different waveforms under the same bandwidth efficiency are simulated and shown in Fig. 6b. AFDM and OTFS are found to be more sensitive to clipping and their BERs are worse than (chirped) DFT-s-OFDM. This is because the signals with high PAPRs, e.g., AFDM and OTFS, are easily affected by clipping. Besides, considering the same bandwidth efficiency using $\mathrm{SF}=4$, OTFS presents higher PAPR than AFDM and that is why its BER degrades more severely than that of AFDM in Fig. 6b.
Figure 6: BERs of different modulation waveforms under the same bandwidth efficiency: a) without clipping and b) with clipping ratio of $1$ [13].
@@ V Conclusion
A new single-carrier waveform, called chirped DFT-s-OFDM, has been proposed for 6G communications, with its performance analyzed and evaluated over delay-Doppler channels with i.i.d. path amplitudes and different integer delays in single-user uplink MAC in this paper. The new waveform by chirping DFT-s-OFDM in the time domain maintains the low PAPR of DFT-s-OFDM. Thanks to the full band transmission and symbols retransmission enabled by chirping and DFT precoding, respectively, the proposed waveform enables enhanced noise suppression of LMMSE equalization. It presents higher output SNR and lower BER than DFT-s-OFDM, AFDM, OTFS, and OFDM in single-user uplink MAC. The PEP analysis and simulation results confirm that the proposed chirped DFT-s-OFDM can achieve full frequency diversity and demonstrate higher resilience to Doppler shifts than DFT-s-OFDM over delay-Doppler channels with i.i.d. path amplitudes and different integer delays.
Note that the proposed chirped DFT-s-OFDM is applicable to multi-user uplink and downlink MAC. Multiple access interference can be mitigated using an ML equalizer to preserve the performance advantage of proposed waveform. In multi-user downlink MAC, the PAPR of (chirped) DFT-s-OFDM would increase but could be still lower than that of AFDM and OFDM. This is because its PAPR is mainly dependent on the value of spreading factor rather than the number of total subcarriers. Since ML equalizer is computationally inefficient and LMMSE equalizer is likely to be susceptible to multiple access interference, future work aims to develop new multi-user receivers for chirped DFT-s-OFDM to eliminate multiple access interference while with low complexity.